# Preferential Attachment Trees with Vertex Death: Persistence of the Maximum Degree

### v1

@@ 4.1. General results
    Suppose there exist $\lambda>0$ and $K\in\mathbb{N}$ such that
    $\widehat{\mu}(\lambda)<1\qquad\text{and}\qquad\prod_{i=K+1}^{\infty}\mathbb{E}\left[\mathrm{e}^{\lambda(X(i)^{\prime}-X(i))}\right]<\infty.$ (PHC)
Remark 4.3.
We observe that the condition that $\sum_{i=1}^{\infty}(X(i)^{\prime}-X(i))$ diverges almost surely in $(a)$, or the condition that $\widehat{\mu}(\lambda)<1$ in (PHC), implies that $\sum_{i=1}^{\infty}X(i)$ diverges almost surely. Indeed, if $\sum_{i=1}^{\infty}X(i)<\infty$ were to hold with positive probability, then
$\bigg{|}\sum_{i=1}^{\infty}(X(i)^{\prime}-X(i))\bigg{|}\leq\sum_{i=1}^{\infty}X(i)^{\prime}+\sum_{i=1}^{\infty}X(i)<\infty,$ (4.1)
with positive probability (where we use that $X(i)^{\prime}\overset{\mathrm{d}}{=}X(i)$ for all $i\in\mathbb{N}$), which contradicts the condition in $(a)$. For the first condition in (PHC),
$\widehat{\mu}(\lambda)=\mathbb{E}\left[\sum_{j=1}^{D}\exp\bigg{(}-\lambda\sum_{i=1}^{j}X(i)\bigg{)}\right]\geq\mathbb{E}\left[D\right]\mathbb{E}\left[\exp\bigg{(}-\lambda\sum_{i=1}^{\infty}X(i)\bigg{)}\right],$ (4.2)
[…]
In the case that the inter-birth times are exponential random variables, we can extend the results of Theorem 4.2 to the existence of $m$-hubs for any $m\in\mathbb{N}$.
Theorem 4.4 (Existence of $m$-hubs).
Let the inter-birth times $(X(i)^{\prime})\overset{\mathrm{d}}{=}(X(i))_{i\in\mathbb{N}}$ and offspring $D$ be defined as after (3.1) such that $X(i)$ follows an exponential distribution for all $i\in\mathbb{N}$ and $\mathbb{P}\!\left(D=\infty\right)>0$. Suppose there exist $\lambda>0$ and $K\in\mathbb{N}$ such that (PHC) is satisfied. Then, for any $m\in\mathbb{N}$, $\mathbb{P}_{\mathcal{S}}$-almost surely $(\mathrm{BP}(t))_{t\geq 0}$ contains a persistent $m$-hub.
The conditions in Theorem 4.2 $(b)$ and Theorem 4.4 ensure the almost sure emergence of a persistent $m$-hub for any $m\in\mathbb{N}$. The following result provides additional conditions under which the persistent $m$-hub is unique, and under which the offspring of the $m$-hubs are distinct.
Theorem 4.5.
Let the inter-birth times $(X(i)^{\prime})\overset{\mathrm{d}}{=}(X(i))_{i\in\mathbb{N}}$ and offspring $D$ be defined as after (3.1) such that Assumption 4.1 and (PHC) are satisfied, and fix $m\in\mathbb{N}$. Suppose that $\mathbb{P}_{\mathcal{S}}$-almost surely $(\mathrm{BP}(t))_{t\geq 0}$ contains a persistent $m$-hub. Also suppose that either
$\sum_{i=1}^{\infty}\mathbb{P}\!\left(X(i)^{\prime}-X(i)\neq 0\right)=\infty\quad\text{or}\quad\mathbb{P}\!\left(X(i)^{\prime}\neq X(i)\right)=1\quad\text{for some }i\in\mathbb{N}.$ (4.3)
Then, $\mathbb{P}_{\mathcal{S}}$-almost surely the process $(\mathrm{BP}(t))_{t\geq 0}$ contains a unique persistent $m$-hub. Additionally, suppose that for any $\varepsilon>0$ we have
$\sum_{i=1}^{\infty}\mathbb{P}\!\left(X(i)>\varepsilon\right)<\infty.$ (4.4)
Then, for any $m\geq 2$ there exists $t^{\prime}>0$ such that $\mathbb{P}_{\mathcal{S}}$-almost surely for all $t>t^{\prime}$,
@@
Remark 4.6.
Theorem 4.5 provides general conditions under which there exists a unique persistent $m$-hub, for any $m$. This requires the existence of a persistent $m$-hub to begin with. For $m=1$ this is true under the same level of generality by Theorem 4.2 $(b)$, whilst we know it to be true for exponentially distributed inter-birth times for $m>1$ by Theorem 4.4. We believe that Theorem 4.4 can be extended to a more general class of inter-birth time distributions as well, for which Theorem 4.5 would then provide the uniqueness of the persistent $m$-hub.
Related work. The results presented in Theorems 4.2, 4.4, and 4.5 are strongly related to recent work of Iyer on persistent hubs in CMJ branching processes [25]. Iyer proves similar results as in Theorem 4.2 and 4.5 (the case $m=1$) under the additional assumption that $D=\infty$, i.e. when all individuals live forever and produce an infinite offspring. Here, based on proof techniques developed by Iyer, we expand this to the case $m>1$ and to the case that $\mathbb{P}\!\left(D=\infty\right)\in(0,1)$. Section 5 deals with Theorem 2.6, which considers the case $\mathbb{P}\!\left(D=\infty\right)=0$.
A further distinction between our results and those of Iyer are the conditions in (4.3), under which we prove that a unique persistent $m$-hub emerges. These are different to the conditions provided by Iyer, as we think that these conditions are too restrictive. Indeed, our proof requires a milder condition under which we can prove that a unique persistent $m$-hub emerges.
4.2. Application to the CTPAVD branching process: Proof of Theorem 2.1
To apply the results in Theorems 4.2, 4.4, and 4.5 to the case of the CTPAVD branching process, we first state the following lemma, which is a slight generalisation of Lemma $3.5$ in [25].
@@
Recall $\widehat{\mu}$ from (3.8). Suppose that $(1)-(3)$ in Assumption 4.1 are satisfied and that there exists $\lambda>0$ such that
$\widehat{\mu}(\lambda)<\infty.$ (4.6)
Then, $|\mathrm{BP}(t)|<\infty$ for all $t\geq 0$ almost surely.
@@
As $\widehat{\mu}(\lambda)$ is decreasing in $\lambda$, we first observe that (4.6), combined with Assumption 4.1, implies that there exists $\lambda^{\prime}\geq\lambda$ such that $\widehat{\mu}(\lambda^{\prime})<1$. Indeed, by using Fubini’s theorem to interchange summation and integration, we can write for any $\alpha\geq 0$,
$\widehat{\mu}(\alpha)=\mathbb{E}\left[\sum_{j=1}^{D}\exp\bigg{(}-\alpha\sum_{i=1}^{j}X(i)\bigg{)}\right]$ $=\sum_{j=1}^{\infty}\mathbb{E}\left[\mathbbm{1}_{\{D\geq j\}}\exp\bigg{(}-\alpha\sum_{i=1}^{j}X(i)\bigg{)}\right]$ (4.7)
$=\sum_{j=1}^{\infty}\mathbb{P}\!\left(D\geq j\right)\mathbb{E}\left[\exp\bigg{(}-\alpha\sum_{i=1}^{j}X(i)\bigg{)}\right],$
where we use the independence of $D$ and the $X(i)$ in the last step. We can split the right-hand side into two terms, namely
$\sum_{j=1}^{\infty}\mathbb{P}\!\left(D\geq j\right)\prod_{i=1}^{j}\mathbb{P}\!\left(X(i)=0\right)+\sum_{j=1}^{\infty}\mathbb{P}\!\left(D\geq j\right)\mathbb{E}\left[\mathbbm{1}_{\{\sum_{i=1}^{j}X(i)>0\}}\exp\bigg{(}-\alpha\sum_{i=1}^{j}X(i)\bigg{)}\right].$ (4.8)
By $(3)$ in Assumption 4.1 there exists $\eta>0$ such that the first sum is at most $1-\eta$. The second sum is finite for any $\alpha\geq\lambda$, since the left-hand side of (4.7) is finite for any $\alpha\geq\lambda$ by assumption. We can thus take $J=J(\eta)\in\mathbb{N}$ sufficiently large, such that for any $\alpha\geq\lambda$,
$\sum_{j=J}^{\infty}\mathbb{P}\!\left(D\geq j\right)\mathbb{E}\left[\mathbbm{1}_{\{\sum_{i=1}^{j}X(i)>0\}}\exp\bigg{(}-\alpha\sum_{i=1}^{j}X(i)\bigg{)}\right]<\frac{\eta}{2}.$ (4.9)
We can then choose $\alpha=\lambda^{\prime}\geq\lambda$ sufficiently large such that
$\sum_{j=1}^{J-1}\mathbb{P}\!\left(D\geq j\right)\mathbb{E}\left[\mathbbm{1}_{\{\sum_{i=1}^{j}X(i)>0\}}\exp\bigg{(}-\lambda^{\prime}\sum_{i=1}^{j}X(i)\bigg{)}\right]<\frac{\eta}{2},$ (4.10)
by the dominated converge theorem (as each summand can be bounded from above by one). Combined, we thus arrive at $\widehat{\mu}(\lambda^{\prime})<1$.
We then prove the desired result. Let $Y$ be an exponential random variable with rate $\lambda^{\prime}$, independent of the offspring random variables $(D^{(v)})_{v\in\mathcal{U}_{\infty}}$ and inter-birth times $(X(vi))_{v\in\mathcal{U}_{\infty},i\in\mathbb{N}}$. Then, for any $v=v_{1}\cdots v_{m}\in\mathcal{U}_{\infty}$, with $m\in\mathbb{N}$ and $v_{1},\ldots,v_{m}\in\mathbb{N}$,
$\mathbb{P}\!\left(\mathcal{B}(v)\leq Y\right)=\mathbb{P}\!\left(D^{(v_{1}\cdots v_{\ell-1})}\geq v_{\ell}\text{ for all }\ell\in[m]\text{, and }\sum_{\ell=1}^{m}\sum_{i=1}^{v_{\ell}}X(v_{1}\cdots v_{\ell-1}i)\leq Y\right).$ (4.11)
We use the independence of the offspring random variables and the inter-birth times, together with the fact that $Y$ is exponentially distributed, to write this as
$\prod_{\ell=1}^{m}\mathbb{P}\!\left(D\geq v_{\ell}\right)\mathbb{E}\left[\exp\bigg{(}-\lambda^{\prime}\sum_{i=1}^{v_{\ell}}X(i)\bigg{)}\right].$ (4.12)
By summing over all individuals $v\in\mathcal{U}_{\infty}$ (where $\mathbb{P}\!\left(\mathcal{B}(\varnothing)\leq Y\right)=1$ as $\mathcal{B}(\varnothing)=0$), we thus obtain
$\mathbb{E}\left[|\mathrm{BP}(Y)|\right]=\sum_{v\in\mathcal{U}_{\infty}}\mathbb{P}\!\left(\mathcal{B}(v)\leq Y\right)=1+\sum_{m=1}^{\infty}\sum_{v_{1}=1}^{\infty}\cdots\sum_{v_{m}=1}^{\infty}\prod_{\ell=1}^{m}\mathbb{P}\!\left(D\geq v_{\ell}\right)\mathbb{E}\left[\exp\bigg{(}-\lambda^{\prime}\sum_{i=1}^{v_{\ell}}X(i)\bigg{)}\right].$ (4.13)
@@
$\sum_{m=0}^{\infty}\Bigg{(}\sum_{j=1}^{\infty}\mathbb{P}\!\left(D\geq j\right)\mathbb{E}\left[\exp\bigg{(}-\lambda^{\prime}\sum_{i=1}^{j}X(i)\bigg{)}\right]\Bigg{)}^{m}=\sum_{m=0}^{\infty}\widehat{\mu}(\lambda^{\prime})^{m},$ (4.14)
where we use (4.7) in the final step. As $\widehat{\mu}(\lambda^{\prime})<1$, the expected value is finite. Hence, $|\mathrm{BP}(Y)|<\infty$ almost surely. Since $Y\geq t$ with positive probability for any $t\geq 0$, it thus follows that $|\mathrm{BP}(t)|<\infty$ for any $t\geq 0$, almost surely. ∎
We then prove Theorem 2.1.
@@
We first check that the CTPAVD branching process satisfies all the conditions in Theorems 4.2, 4.4, and 4.5. We then translate this back to the PAVD model.
For the CTPAVD branching process, we have
$X(i)\overset{\mathrm{d}}{=}X(i)^{\prime}\overset{\mathrm{d}}{=}\text{Exp}(b(i-1)+d(i-1))\qquad\text{for all }i\in\mathbb{N},$ (4.15)
and $D$ is as in (2.6). It directly follows that $(1)-(3)$ of Assumption 4.1 are satisfied, and $(4)$ of Assumption 4.1 is equivalent to Assumption (F-D) not being satisfied, due to Lemma 5.1 (see also [23, Lemma $5.3$]).
For $(a)$ in Theorem 4.2 it remains to check that $\sum_{i=1}^{\infty}(X(i)^{\prime}-X(i))$ diverges almost surely. As the summands are independent and symmetric, we have
${\rm Var}\bigg{(}\sum_{i=1}^{n}(X(i)^{\prime}-X(i))\bigg{)}=2\sum_{i=1}^{n}{\rm Var}(X(i))=\sum_{i=1}^{n}\frac{2}{(b(i)+d(i))^{2}},$ (4.16)
which has a finite limit as $n$ tends to infinity if and only if Assumption (C-V) is not satisfied. Kolmogorov’s two and three series theorems (see e.g. [19, Theorems $2.5.3$ and $2.5.4$]) thus imply that $\sum_{i=1}^{\infty}(X(i)^{\prime}-X(i))$ converges almost surely if and only if Assumption (C-V) is not satisfied, so that $(a)$ follows. So, $\mathbb{P}_{\mathcal{S}}$-almost surely $(\mathrm{BP}(t))_{t\geq 0}$ does not contain a persistent $m$-hub, for any $m\in\mathbb{N}$, when Assumption (C-V) is not satisfied.
For $(b)$, we directly conclude from the above that $\sum_{i=1}^{\infty}(X(i)^{\prime}-X(i))$ converges when Assumption (C-V) is satisfied. It remains to check the conditions in (PHC). For the second condition in (PHC), we first observe that $b+d$ tends to infinity when Assumption (C-V) is satisfied. For any $\lambda>0$ we can thus choose $K\in\mathbb{N}$ sufficiently large such that $b(i)+d(i)\geq\lambda\sqrt{2}$ for all $i\geq K$. Now,
$\prod_{i=K+1}^{\infty}\mathbb{E}\left[\mathrm{e}^{\lambda((X(i)^{\prime}-X(i))}\right]=\prod_{i=K+1}^{\infty}\frac{(b(i)+d(i))^{2}}{(b(i)+d(i)+\lambda)(b(i)+d(i)-\lambda)}=\prod_{i=K+1}^{\infty}\frac{(b(i)+d(i))^{2}}{(b(i)+d(i))^{2}-\lambda^{2}}.$ (4.17)
Taking the logarithm of the right-hand side yields
$-\sum_{i=K+1}^{\infty}\log\Big{(}1-\frac{\lambda^{2}}{(b(i)+d(i))^{2}}\Big{)}.$ (4.18)
As $b(i)+d(i)\geq\lambda\sqrt{2}$ for all $i\geq K$ and since $\log(1-x)\geq-x-x^{2}$ for $x\in(0,1/2)$, we thus arrive at the upper bound
$\lambda^{2}\sum_{i=K+1}^{\infty}\frac{1}{(b(i)+d(i))^{2}}+\lambda^{4}\sum_{i=K+1}^{\infty}\frac{1}{(b(i)+d(i))^{4}}\leq\frac{3}{2}\lambda^{2}\sum_{i=K+1}^{\infty}\frac{1}{(b(i)+d(i))^{2}}$ (4.19)
which is finite for any $\lambda>0$ when Assumption (C-V) is satisfied. The first condition in (PHC) is implied by Assumption (F-L), i.e. that $\widehat{\mu}(\lambda)<\infty$ for some $\lambda>0$. We then, as in the first part of the proof of Lemma 4.7, conclude that there exists $\lambda^{\prime}\geq\lambda$ such that $\widehat{\mu}(\lambda^{\prime})<1$. We note that the second condition in (PHC) is satisfied for all $\lambda>0$ and any $K\geq\lambda\sqrt{2}$, so that is satisfied for $\lambda^{\prime}$ and $K\geq\lambda^{\prime}\sqrt{2}$ in particular. So, with $\lambda^{\prime}$ and $K\geq\lambda^{\prime}\sqrt{2}$, we conclude that (PHC) is satisfied, so that part $(b)$ of Theorem 4.2 holds. Hence, for any $m\in\mathbb{N}$, $\mathbb{P}_{\mathcal{S}}$-almost surely $(\mathrm{BP}(t))_{t\geq 0}$ contains a persistent $m$-hub when Assumptions (C-V) and (F-L) are satisfied. Theorem 4.4 then also directly follows.
To obtain Theorem 4.5, we note that the exponential distribution has no atoms, so that there is a unique persistent $m$-hub $\mathbb{P}_{\mathcal{S}}$-almost surely, for any $m\in\mathbb{N}$. Then,
$\mathbb{P}\!\left(X(i)>\varepsilon\right)=\mathrm{e}^{-\varepsilon(b(i)+d(i))}\leq\frac{1}{(b(i)+d(i))^{2}},$ (4.20)
where the inequality follows for any $\varepsilon>0$ and all $i\geq I_{\varepsilon}$, for some fixed $I_{\varepsilon}\in\mathbb{N}$, from the fact that $b(i)+d(i)$ tends to infinity with $i$ when Assumption (C-V) holds. As a result, the second condition in (4.4) is satisfied (again applying Assumption (C-V)), so that, for any $m\in\mathbb{N}$, $\mathbb{P}_{\mathcal{S}}$-almost surely the offspring of the persistent $1$-hub, persistent $2$-hub, through the persistent $m$-hub are distinct for all $t>t^{\prime}$.
To summarise, we have the following table when the inter-birth times satisfy (4.15) and $D$ is as in (2.6).
Assumptions Result
(C-V) and (F-D) not satisfied Theorem 4.2 (a): No persistent $m$-hub
(C-V) and (F-L) satisfied, (F-D) not satisfied Theorems 4.2 (b) and 4.5: Unique persistent $m$-hub with distinct degrees
Finally, we translate this result back to the PAVD model, for which we use the embedding in Proposition 3.1. First, we claim that $\lim_{n\to\infty}\tau_{n}=\infty$ $\mathbb{P}_{\mathcal{S}}$-almost surely. This follows since $\mathbb{P}\!\left(\mathcal{S}\right)>0$ and since $\widehat{\mu}(\lambda)<\infty$, so that Lemma 4.7 yields that $\mathrm{BP}(t)$ is finite for all $t\geq 0$, almost surely. Hence, $\lim_{n\to\infty}\tau_{n}=\infty$ $\mathbb{P}_{\mathcal{S}}$-almost surely. As a result, for any $t^{\prime}>0$ there exists $n_{0}\in\mathbb{N}$ such that $\tau_{n}>\tau_{n_{0}}\geq t^{\prime}$ for all $n\geq n_{0}$ $\mathbb{P}_{\mathcal{S}}$-almost surely. Under the conditions in Theorem 2.1 it thus follows that $(T_{n})_{n\in\mathbb{N}}$ does not contain a persistent $m$-hub $\mathbb{P}_{\mathcal{S}}$-almost surely for any $m\in\mathbb{N}$ when Assumption (C-V) is not satisfied. When Assumption (C-V) is satisfied and Assumption (F-L) also holds, then for any $m\in\mathbb{N}$ there exists $t^{\prime}>0$ and random variables $I^{(1),\mathrm{ct}},\ldots,I^{(m),\mathrm{ct}}$, such that
$(I^{(1),\mathrm{ct}}_{t},\ldots,I^{(m),\mathrm{ct}}_{t})=(I^{(1),\mathrm{ct}},\ldots,I^{(m),\mathrm{ct}})$ (4.21)
for all $t>t^{\prime}$ $\mathbb{P}_{\mathcal{S}}$-almost surely. As a result, there exists $n_{0}\in\mathbb{N}$ and random variables $I^{(1)},\ldots,I^{(m)}$, such that
$(I_{n}^{(1)},\ldots,I_{n}^{(m)})=(I^{(1)},\ldots,I^{(m)})$ (4.22)
for all $n\geq n_{0}$ $\mathbb{P}_{\mathcal{S}}$-almost surely. Equivalently, $(T_{n})_{n\in\mathbb{N}}$ contains a persistent $m$-hub $\mathbb{P}_{\mathcal{S}}$-almost surely, for any $m\in\mathbb{N}$, when Assumptions (C-V) and (F-L) are satisfied.
Since the exponential distribution has no atoms, it follows from the first part of Theorem 4.5 that when Assumption (C-V) is satisfied, for any $m\in\mathbb{N}$, $\mathbb{P}_{\mathcal{S}}$-almost surely $(\mathrm{BP}(t))_{t\geq 0}$ contains a unique persistent $m$-hub. It then directly follows that $(T_{n})_{n\in\mathbb{N}}$ contains a unique persistent $m$-hub $\mathbb{P}_{\mathcal{S}}$-almost surely as well, for any $m\in\mathbb{N}$. We remark that this also directly follows from the observation that at most one vertex increases its degree at every step in the discrete tree process $(T_{n})_{n\in\mathbb{N}}$, so that at most one vertex can attain the $m^{\mathrm{th}}$ largest degree for all but finitely many steps.
Finally, for any $m\in\mathbb{N}$ there exists $t^{\prime}>0$ and random variables $I^{(1),\mathrm{ct}},\ldots,I^{(m),\mathrm{ct}}$ such that
$(I^{(1),\mathrm{ct}}_{t},\ldots,I^{(m),\mathrm{ct}}_{t})=(I^{(1),\mathrm{ct}},\ldots,I^{(m),\mathrm{ct}})$ (4.23)
for all $t>t^{\prime}$ $\mathbb{P}_{\mathcal{S}}$-almost surely. Then, there exists $t^{\prime\prime}>t^{\prime}$ such that
$\deg_{t}(I^{(1),\mathrm{ct}})>\cdots>\deg_{t}(I^{(m),\mathrm{ct}})$ (4.24)
for all $t>t^{\prime\prime}$ $\mathbb{P}_{\mathcal{S}}$-almost surely. As $\tau_{n}$ tends to infinity with $n$ $\mathbb{P}_{\mathcal{S}}$-almost surely, and
$\{\deg_{n}(i):i\in\mathcal{A}_{n}\}\overset{\mathrm{d}}{=}\{\deg_{\tau_{n}}(v):v\in\mathcal{A}_{\tau_{n}}^{\mathrm{ct}}\}$ (4.25)
by Proposition 3.1, we thus obtain that there exists $n_{1}\in\mathbb{N}$ such that
$\deg_{n}(I^{(1)})>\cdots>\deg_{n}(I^{(m)})$ (4.26)
for all $n\geq n_{1}$ $\mathbb{P}_{\mathcal{S}}$-almost surely, which together with (4.22) yields (2.12) and concludes the proof. ∎
The remainder of this section is devoted to proving Theorems 4.2, 4.4, and 4.5.

### v2

@@ 3. Methodology: Embedding in CMJ branching process
    If Assumption (C-V) is satisfied and Assumption (F-L) is satisfied, then for any $m\in\mathbb{N}$, $\mathbb{P}_{\mathcal{S}}$-almost surely $(\mathrm{BP}(t))_{t\geq 0}$ contains a unique persistent $m$-hub, in the sense of (3.18). For any $m\geq 2$ there exists $t^{\prime}>0$ such that $\mathbb{P}_{\mathcal{S}}$-almost surely for all $t>t^{\prime}$,
@@
Related work. The results presented in Theorem 4.1 are related to recent work of Iyer on persistent hubs in general CMJ branching processes [25]. The main difference between the results of Iyer and the results presented here are:
  • Iyer focusses on general branching processes, where the inter-birth times $X(i)$ are not necessarily exponentially distributed.
  • Iyer focusses the case where the offspring $D^{(v)}$ is infinite for each individual $v$.
  • The results of Iyer are restricted to the case $m=1$ and do not include (4.1).
In short, based on proof techniques developed by Iyer, we extend his results to the case $m>1$ and to the case that $\mathbb{P}\!\left(D=\infty\right)\in(0,1]$ (which is the case when Assumption (F-D) is not satisfied) when the inter-birth times are exponentially distributed. These extensions can be proved for more general CMJ branching processes, where we do not assume the inter-birth times to be exponentially distributed, as well. As the focus of this work lies on the PAVD model, we do not prove this explicitly, but rather provide a discussion on how to extend (part of) the results in Theorem 4.1 to general CMJ branching processes can be found Appendix A.
4.1. Proof of Theorem 2.1
To translate the results in Theorem 4.1 to the results of Theorem 2.1, we first state the following lemma, which is a slight generalisation of Lemma $3.5$ in [25].
@@
Recall the Laplace transform $\widehat{\mu}$ from (2.8) and suppose that Assumption (F-L) is satisfied. That is, suppose that there exists $\lambda>0$ such that $\widehat{\mu}(\lambda)<\infty$. Then, $|\mathrm{BP}(t)|<\infty$ for all $t\geq 0$ almost surely.
@@
We start by observing that Assumption (F-L) is equivalent to there existing $\lambda^{\prime}\geq\lambda$ such that $\widehat{\mu}(\lambda^{\prime})<1$. Indeed, since $\widehat{\mu}(\lambda)<\infty$ and $\widehat{\mu}(\alpha)$ is decreasing in $\alpha$, it follows from the dominated convergence theorem that
$\lim_{\alpha\to\infty}\widehat{\mu}(\alpha)=\mu(\{0\})=0\qquad\text{and so there exists $\lambda^{\prime}\in[\lambda,\infty)$ such that }\widehat{\mu}(\lambda^{\prime})<1.$ (4.2)
We then prove the result. For any $v=v_{1}\cdots v_{m}\in\mathcal{U}_{\infty}$, with $m\in\mathbb{N}$ and $v_{1},\ldots,v_{m}\in\mathbb{N}$,
$\mathbb{P}\!\left(\mathcal{B}(v)\leq t\right)=\mathbb{P}\!\left(D^{(v_{1}\cdots v_{\ell-1})}\geq v_{\ell}\text{ for all }\ell\in[m]\text{, and }\sum_{\ell=1}^{m}\sum_{i=0}^{v_{\ell}-1}E^{(v_{1}\cdots v_{\ell-1})}_{i}\leq t\right).$ (4.3)
We use the independence of the offspring random variables and the inter-birth times, together with a Chernoff bound with $\lambda^{\prime}$, to bound the right-hand side from above by
$\mathrm{e}^{\lambda^{\prime}t}\prod_{\ell=1}^{m}\mathbb{P}\!\left(D\geq v_{\ell}\right)\mathbb{E}\left[\exp\bigg(-\lambda^{\prime}\sum_{i=0}^{v_{\ell}-1}E_{i}\bigg)\right].$ (4.4)
By summing over all individuals $v\in\mathcal{U}_{\infty}$ (where $\mathbb{P}\!\left(\mathcal{B}(\varnothing)\leq t\right)=1$ as $\mathcal{B}(\varnothing)=0$), we obtain
$\mathbb{E}\left[|\mathrm{BP}(t)|\right]$ $=\sum_{v\in\mathcal{U}_{\infty}}\mathbb{P}\!\left(\mathcal{B}(v)\leq t\right)$ (4.5)
$\leq 1+\mathrm{e}^{\lambda^{\prime}t}\sum_{m=1}^{\infty}\sum_{v_{1}=1}^{\infty}\cdots\sum_{v_{m}=1}^{\infty}\prod_{\ell=1}^{m}\mathbb{P}\!\left(D\geq v_{\ell}\right)\mathbb{E}\left[\exp\bigg(-\lambda^{\prime}\sum_{i=0}^{v_{\ell}-1}E_{i}\bigg)\right].$
@@
$\sum_{m=0}^{\infty}\Bigg(\sum_{j=1}^{\infty}\mathbb{P}\!\left(D\geq j\right)\mathbb{E}\left[\exp\bigg(-\lambda^{\prime}\sum_{i=0}^{j-1}E_{i}\bigg)\right]\Bigg)^{m}.$ (4.6)
By rewriting the inner sum and using that $D$ is independent of $(E_{i})_{i\in\mathbb{N}}$, we obtain
$\sum_{j=1}^{\infty}\mathbb{P}\!\left(D\geq j\right)\mathbb{E}\left[\exp\bigg(-\lambda^{\prime}\sum_{i=0}^{j-1}E_{i}\bigg)\right]=\mathbb{E}\left[\sum_{j=1}^{D}\exp\bigg(-\lambda^{\prime}\sum_{i=0}^{j-1}E_{i}\bigg)\right]=\widehat{\mu}(\lambda^{\prime}),$ (4.7)
where we recall $\widehat{\mu}$ from (2.8). We finally obtain that $\mathbb{E}\left[|\mathrm{BP}(t)|\right]\leq\mathrm{e}^{\lambda^{\prime}t}\sum_{m=0}^{\infty}\widehat{\mu}(\lambda^{\prime})^{m}$, which is finite since $\widehat{\mu}(\lambda^{\prime})<1$. Hence, $|\mathrm{BP}(t)|<\infty$ almost surely, as desired. ∎
Theorem 2.1 then readily follows.
@@
Recall the stopping times $(\tau_{n})_{n\in\mathbb{N}}$ from (3.11). By using Lemma 4.2 we observe that $\tau_{n}\to\infty$ almost surely conditionally on survival. As a result, by using Proposition 3.1 we obtain that Theorem 4.1 directly implies Theorem 2.1. ∎
The remainder of this section is devoted to proving Theorem 4.1.